OpenStudy (anonymous):

the length of a rectangle is 6 more than the width, the area of a rectangle is 16m^2 find the length and width

OpenStudy (anonymous):

Let \(x\) be the width and \(y\) be the length. What sort of equations do they tell us?

OpenStudy (anonymous):

"the length of a rectangle is 6 more than the width" How to we write this in terms of \(x\) and \(y\)?

OpenStudy (anonymous):

"the area of a rectangle is 16m^2 " and how do we write this in algebra?

OpenStudy (anonymous):

@ricque Got an idea?

OpenStudy (anonymous):

whats that

OpenStudy (anonymous):

Here's a hint: "the length of a rectangle is 6 more than the width" becomes "y is 6 more than x" since we said "Let x be the width and y be the length." Can you

OpenStudy (anonymous):

Can you make it into a full on math equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1) Let x be the width and y be the length. (our choice) 2) "the length of a rectangle is 6 more than the width" "y is 6 more than x" "y = 6 more than x" y = 6 + x 3) "the area of a rectangle is 16m^2 " "xy is 16m^2" xy = 16m^2 So we have two equations: y = 6 + x xy = 16 Can you solve it?